40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 3 01:46:54 2011, Kevin, Update, Electronics, POY Shot Noise and Dark Spectrum 
    Reply  Thu Feb 3 11:13:52 2011, Koji, Update, Electronics, POY Shot Noise and Dark Spectrum 
Message ID: 4242     Entry time: Thu Feb 3 01:46:54 2011     Reply to this: 4244
Author: Kevin 
Type: Update 
Category: Electronics 
Subject: POY Shot Noise and Dark Spectrum 

[Koji and Kevin]

I measured the shot noise of POY and fit the data to determine the RF transimpedance at 11 MHz and the dark current. The transimpedance is (3.860 +- 0.006) kΩ. I realize that there are not many data points past the dark current but I did not want to take any further data because the light bulb was getting pretty bright. If this is a problem, I can try to redo the measurement using a lens to try to focus more of the light from the bulb onto the photodiode.

I also measured the spectrum and recorded a time series of the RF signal with the light to the photodiode blocked. These measurements do not show any large oscillations like the ones found for POX.

The plots of the measurements are on the wiki at http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/POY.

ELOG V3.1.3-